the measure of two complementary angles are 2x-10 and 4x-2

Answer by LinnW(1048)

(2x-10) + (4x-2) = 90 , two complementary angles total 90 degrees
6x -12 = 90
add 12 to each side
6x = 102
x = 17
The first angle (2x-10) = 2(17) - 10 = 34 -10 = 24
The second angle (4x-2) = 4(17) - 2 = 68 -2 = 66